OpenStudy (anonymous):

I need help with problems involving the square root property. (5x-1)^2=4/25

OpenStudy (anonymous):

Get rid of the squared on both sides of the equation by square rooting both sides. - this will give you 5x-1= +2/5 or 5x-1 = -2/5 Add 1 to both sides of the equation. - This will give you 5x= 7/5 or 5x= 3/5 Divide both sides of the equation by 5 (or I usually multiply by the inverse; 1/5 in this case). - This will give you the answer: x= 7/25 or x=3/25

OpenStudy (anonymous):

Thanks! That really helped.